Write an equation in slope-intercept form of the line that passes through the given point and is perpendicular to the graph of the given equation. (-3,5); y=1/3x-3
Answer by stanbon(75887)

Write an equation in slope-intercept form of the line that passes through the given point and is perpendicular to the graph of the given equation. (-3,5); y=1/3x-3
------
The given equation has slope = 1/3
----
A perpendicular line must have slope = -3
-----
Form:: y = mx + b
Substitute and solve for "b":
5 = -3*-3 +b
b = -4
------
Ans: y = -3x -4
=================
Cheers,
Stan H.
================